Use the References to access important values if needed for this question.



What volume of a 0.226 M hydrochloric acid solution is required to neutralize 11.8 mL of a 0.191 M sodium hydroxide solution?


_________ mL hydrochloric acid

1
Expert's answer
2021-10-12T02:07:11-0400

HCl+NaOH→H2O+NaCl

Molar mass of sodium hydroxide = 39.997 g/mol

39.997×0.191= 7.6394

11.8÷7.6394= 1.545mol

Molar mass of HCl = 36.458 g/mol

1.545× 36.458= 56.33g

36.458× 0.226= 8.24g

56.33÷8.24= 6.84ml